I can not set up the two linear equations for the following word problem.
THERE ARE TWO FLOOR PLANS FOR NEW HOMES. Floor plan #1 and Floor plan #2 and there are 56 houses avalible.
There are three times as many floor plan #2 as there are floor plan #1.
I tried
X+Y =56
x+3y=56
But mathmatically it is not comming out right. I also know the answer is 14 #1 floor plans and 42 #2 floor plans but I am having a problem setting up the equations.

Answer by checkley77(12844)

X+Y=56
X=3Y
3Y+Y=56
4Y=56
Y=56/4
Y=14 FOR #1
56-14=42 FOR #2
PROOF:4*3=42